OpenStudy (anonymous):

-4(7k+3)=27 what does k equal

OpenStudy (anonymous):

-4*(7k+3)=27 -4*(7k+3)/-4=27/-4 7k+3=-6.75 7k+3-3=-6.75-3 7k = -9.75 7k/7=-9.75/7 k= -1.393

OpenStudy (anonymous):

i think... its probably sketchy...

OpenStudy (anonymous):

its ok thanks

OpenStudy (anonymous):

ur welcome :)